OpenStudy (vortish):

use the symbols to right the expression and then evaluate it sixteen squared divided by four cubed i think its 16^2 by 4^3 am I wrong

OpenStudy (anonymous):

I believe that you have it right. \[\frac{16^{2}}{4^{3}}\]Do you know how to evaluate it?

OpenStudy (vortish):

i think so but i could be wrong i think its 36/64

OpenStudy (anonymous):

\[\frac{16^{2}}{4^{3}} = \frac{4^{2(2)}}{4^{3}} = \frac{4^{4}}{4^{3}} = 4^{4 - 3} = 4^{1} = 4\]

OpenStudy (vortish):

do i need to write it out that way calcmathlete

OpenStudy (anonymous):

Well, I showed you how to do it. You don't have to write it that way so long as you know how to get the answer which is 4.

OpenStudy (vortish):

thanks

OpenStudy (anonymous):

Glad to help :)